A triangle has sides with lengths in centimeters that are consecutive even integers. Find the lengths if the perimeter is 186 cm.
Answer by Theo(2978)

x = first side
x + 2 = second side
x + 4 = third side
-----
sum of 3 sides equals 3x + 6
-----
3x + 6 = 186
3x = 180
x = 60
x+2 = 62
x+4 = 64
-----
60 + 62 + 64 = 186
-----